# Can LLMs Predict Citation Intent? An Experimental Analysis of In-context Learning and Fine-tuning on Open LLMs

### v2

@@ Abstract
1 Introduction
Citations are references of research articles to external sources of information included to support claims, provide context, criticize, or acknowledge prior work. Although their primary function is to inform and redirect the reader, citations are also frequently used for other purposes, such as serving as proxies of scientific impact in various types of analysis [18]. In such cases, understanding the exact intent of a citation is crucial, as not all types of citations should be considered. For instance, while measuring the impact of an article, citations that criticize the work should generally not be considered as contributing to its impact.
Predicting citation intent based on its context (i.e., the sentences in the manuscript that accompany the citation) and other related information, has become an important classification problem to support the aforementioned use cases. Existing approaches have traditionally relied on linguistic features [17], machine learning methods [32], and, in recent years, on domain-specific pre-trained language models (PLMs), such as SciBERT [3], that require large scientific datasets such as [21] and task-specific architectures111In this paper, when referring to traditional PLMs in this context, we primarily mean models based on the encoder-only Transformer architecture, such as BERT and its derivatives (e.g., SciBERT), which were commonly fine-tuned for specific classification tasks. This distinguishes them from the general-purpose Large Language Models (LLMs) evaluated in this work, which typically utilize decoder-only or encoder-decoder Transformer architectures and are often leveraged for their generative and in-context learning capabilities.. In contrast, this work is the first to explore the potential of open, general-purpose Large Language Models (LLMs) – which, despite potentially encountering scientific text during their broad pre-training, are not specifically optimized for the scientific domain like SciBERT – to accurately identify citation intent, evaluating their effectiveness through in-context learning and fine-tuning on minimal task-specific data.
@@ 1 Introduction
In-context learning is a paradigm that allows LLMs to learn tasks given only a few examples in the form of demonstration [4, 8]. In-context learning is particularly suited to citation intent classification, as it allows the models to leverage their contextual understanding of language to make accurate predictions.
@@ 2.1 Models
For our experiments, we selected the Instruction-tuned versions of five prominent open-weight model families: LLaMA [11], Mistral [15], Phi [1], Gemma [10], and Qwen [2, 26]. A key challenge was performing the experiments on commodity hardware, both due to computational limitations and to demonstrate that citation intent classification can be executed efficiently with limited resources. This influenced the selection of models, which range in size from small (B parameters) to medium-sized (B parameters).
@@
To reduce the memory footprint and computational requirements of our evaluation, we utilize the 8-bit (Q8) quantized versions of the models. This approach significantly reduces memory usage without compromising performance. Quantization involves converting model parameters from higher numerical precisions (e.g. 16-bit floating-point) to formats of lower numerical precision (e.g. 8-bit integers). This process enables more efficient computation and reduced memory usage while largely preserving model performance and expressive power [16].
@@ 2.3 Configuration Parameters
To evaluate model performance, we applied four prompting methods444Throughout the paper, we use “In-context Learning Methods” and “Prompting Methods” interchangibly. by following the In-Context Learning paradigm [4]: Zero-shot (no examples), One-shot (a single example per class), Few-shot ( examples per class), and Many-shot ( examples per class).
These prompting methods were selected to evaluate whether model performance improves as the number of examples increases, and to identify any saturation point where additional examples yield diminishing or negative returns. This aligns with prior literature on in-context learning methods, where these specific configurations have been extensively studied [8]. Incorporating examples directly into the prompts allows models to better understand the task and the expected output format, which is especially important for citation intent classification, where the citation context is key to determining the correct label.
@@
SP2: This prompt introduced more structure, drawing inspiration from the CO-STAR framework [31]. We adapted relevant CO-STAR components by definint the task Context, the Objective of the model, and detailed Response Rules. Given that our task requires a single, specific class label output, the stylistic components of CO-STAR (Style, Tone, and Audience) were not incorporated.
@@
Inline: Example citation sentences and their corresponding classes (structured according to the defined query templates) are embedded directly and sequentially within the main prompt, immediately preceding the actual query sentence that the model needs to classify.
@@
Temperature is a hyperparameter that controls the randomness or creativity of a language model’s outputs by adjusting the probability distribution of possible next tokens [16]. Lower temperatures (i.e., close to ) correspond to greedy decoding, where the model deterministically selects the most probable token at each step, while higher temperatures (closer to ) introduce greater variability by allowing the model to sample from a broader range of options.
@@ 3 Experimental Results and Analysis
This section presents the results of our experiments, focusing on model performance across various configurations and parameter settings. An initial overview of the peak F1-score achieved by each of the twelve models on both the SciCite and ACL-ARC datasets is summarized in Table 1. The subsections that follow provide a detailed analysis of top-performing model configurations, evaluate overall model rankings, and explore the influence of specific parameters.
SciCite | ACL-ARC | ||||
Rank | Model | F1-Score | Rank | Model | F1-Score |
1 | Qwen 2.5 – 14B | 78.33 | 1 | Qwen 2.5 – 14B | 63.68 |
2 | Gemma 2 – 27B | 77.86 | 2 | Gemma 2 – 27B | 58.95 |
3 | Mistral Nemo – 12B | 77.39 | 3 | Gemma 2 – 9B | 57.19 |
4 | Gemma 2 – 9B | 75.12 | 4 | Qwen 2 – 7B | 51.26 |
5 | Phi 3 Medium – 14B | 74.67 | 5 | LLaMA 3.1 – 8B | 48.45 |
6 | LLaMA 3 – 8B | 74.39 | 6 | Mistral Nemo – 12B | 48.11 |
7 | Qwen 2 – 7B | 72.89 | 7 | Phi 3.5 Mini – 3.8B | 43.74 |
8 | LLaMA 3.1 – 8B | 72.46 | 8 | Phi 3 Medium - 14B | 43.46 |
9 | Gemma 2 - 2B | 68.79 | 9 | Gemma 2 - 2B | 40.96 |
10 | Phi 3.5 Mini – 3.8B | 68.25 | 10 | LLaMA 3.2 - 3B | 40.07 |
11 | LLaMA 3.2 - 3B | 67.99 | 11 | LLaMA 3 - 8B | 38.06 |
12 | LLaMA 3.2 - 1B | 45.44 | 12 | LLaMA 3.2 - 1B | 24.60 |
@@ 3.1 Model Performance over Configurations
SciCite | ACL-ARC | ||||
Rank | Model | Ranking Score | Rank | Model | Ranking Score |
1 | Qwen 2.5 – 14B | 144.163 | 1 | Qwen 2.5 – 14B | 146.496 |
2 | Mistral Nemo – 12B | 68.479 | 2 | Gemma 2 – 27B | 62.973 |
3 | Gemma 2 – 27B | 63.369 | 3 | Gemma 2 – 9B | 55.318 |
4 | Gemma 2 – 9B | 59.760 | 4 | Qwen 2 – 7B | 41.496 |
5 | Qwen 2 – 7B | 38.617 | 5 | LLaMA 3.1 – 8B | 40.844 |
6 | LLaMA 3.1 – 8B | 27.624 | 6 | Phi 3.5 Mini – 3.8B | 30.470 |
7 | LLaMA 3 – 8B | 26.159 | 7 | Mistral Nemo – 12B | 30.285 |
8 | Phi 3.5 Mini – 3.8B | 22.088 | 8 | LLaMA 3.2 - 3B | 23.984 |
9 | Phi 3 Medium – 14B | 19.370 | 9 | Phi 3 Medium - 14B | 21.269 |
10 | Gemma 2 - 2B | 19.253 | 10 | LLaMA 3.2 - 1B | 16.796 |
11 | LLaMA 3.2 - 3B | 18.286 | 11 | Gemma 2 - 2B | 16.534 |
12 | LLaMA 3.2 - 1B | 14.136 | 12 | LLaMA 3 - 8B | 16.359 |
3.1.2 Ranked Evaluation.
While the Best-Performer Evaluation provided insights into the most dominant models across configurations, it lacked the ability to capture nuances among high-performing models. For instance, a model that consistently ranked second in multiple configurations would not be reflected in that approach. To address this limitation, we adopted a ranking methodology inspired by Reciprocal Rank Fusion (RRF) [6]. This approach allowed us to evaluate all models considering their relative performance across different configurations.
@@
Table 3 presents the results of this ranked evaluation. The ranked evaluation aligns with the Best-Performer Evaluation in identifying Qwen 2.5 14B as the most dominant model. This view also allows us to gain insights into the relative performance of other the models, highlighting distinctions among high performers, as well as the performance of lower-ranked models.
@@ 3.2 Parameter Performance Analysis
3.2.1 Chi-Square Statistical Test.
To complement our parameter performance analysis, we also conducted a Chi-Square Test of Independence [12] to evaluate the relationship between parameter settings and F1-scores. This approach allowed us to determine whether specific parameter settings were significantly associated with performance outcomes and validate the results of our primary analysis. The results of the test (summarized in Table 5) largely reinforces the findings from our quantile-based analysis regarding the influence of different parameter settings.
@@
4 Fine-tuning
4.1 Training Configuration
For this experiment, we used the SciCite and ACL-ARC datasets, which were converted into the Alpaca format [30]. This format includes a system prompt, an instruction, the citing sentence, and the true label. The use of Supervised Fine-Tuning (SFT) was motivated by its ability to adapt pre-trained models to specific tasks using minimal labeled data, making it an effective approach for citation intent classification.
To fine-tune the Qwen 2.5 14B Instruct model, we used LLaMA-Factory [33] on an AWS g6e.12xlarge EC2 instance equipped with 4 NVIDIA L40S GPUs, providing a combined GPU memory of 192GB. The training process was conducted using fp16 mixed precision to optimize memory usage and speed. To prevent memory issues during fine-tuning, we employed DeepSpeed ZeRO Stage 3 Offload [27], which enabled efficient memory management by offloading optimizer states and gradients to the CPU. The training parameters included a learning rate of 5e-5, a batch size of 16, and 10 epochs. The model was optimized using AdamW, with a cutoff length of 512 tokens for input sequences.
To optimize the fine-tuning process, we employed Low-Rank Adaptation (LoRA), a widely used parameter-efficient fine-tuning (PEFT) method [14, 13]. LoRA enables efficient adaptation of large language models by freezing the pre-trained model weights and introducing trainable low-rank matrices into the Transformer layers, significantly reducing the number of trainable parameters while maintaining performance. For this experiment, we configured LoRA with a rank of 8, alpha of 16, and a dropout of 0.1, which allowed us to fine-tune the model effectively without exceeding memory constraints. The outlined training parameters are also summarized in Table 6.
@@ 4.2 Results
Model | SciCite | ACL-ARC |
Qwen 2.5 – 14B Instruct Q8 | 78.33 | 63.68 |
Qwen 2.5 – 14B Instruct FP16 | 80.41 | 65.64 |
Qwen 2.5 – 14B Fine-tuned Q8 | 86.47 | 68.48 |
Qwen 2.5 – 14B Fine-tuned FP16 | 86.84 | 67.73 |
@@ 4.3 Discussion
We consider a key advantage of LLMs to be their ease of use and deployment. Tools such as LM Studio and Ollama888https://ollama.com/ allow models like those outlined in this paper to be deployed locally with zero technical expertise, making them accessible to users without a computer science background. In addition to their accessibility, LLMs offer significant adaptability. Unlike traditional methods, which often require complex pretraining or domain-specific tuning, LLMs can be fine-tuned for a wide range of scientometric tasks, such as citation recommendation, paper summarization, or trend analysis, without the need for bespoke architectures. Furthermore, LLMs can scale effectively to new scientific domains or datasets, requiring only small amounts of task-specific data to adapt to underexplored scientific fields. , our fine-tuned models required only the several thousand citing sentences provided by our task-specific datasets for adaptation. This contrasts sharply with models like SciBERT, whose effectiveness stems from deliberate pre-training exclusively on a large corpus of scientific papers (millions of articles), optimizing it for scientific language understanding. While the general-purpose LLMs used in our study were likely exposed to scientific text within their vast, diverse pre-training data drawn from the web, this exposure is incidental rather than targeted. They were not specifically pre-trained or architected with the primary goal of processing scientific literature, unlike SciBERT. Our results demonstrate that even without such specialized scientific pre-training, general-purpose LLMs can achieve competitive performance through efficient fine-tuning on minimal task-specific data.
The promising results observed in this study suggest that the performance of LLMs in citation intent classification can be further improved with techniques such as chain-of-thought prompting and reasoning-focused models. These methods could enhance the models’ ability to better distinguish subtle differences in citation intent, refining their predictions and improving overall classification accuracy.

### v3

@@ Abstract
Can LLMs Predict Citation Intent? An Experimental Analysis of In-context Learning and Fine-tuning on Open LLMs
Paris Koloveas IMSI, ATHENA RC / Athens, GR University of the Peloponnese / Tripolis, GR pkoloveas@athenarc.gr Serafeim Chatzopoulos IMSI, ATHENA RC / Athens, GR schatz@athenarc.gr
Thanasis Vergoulis IMSI, ATHENA RC / Athens, GR vergoulis@athenarc.gr Christos Tryfonopoulos University of the Peloponnese / Tripolis, GR trifon@uop.gr
Citations are references of research articles to external sources of information included to support claims, provide context, criticize, or acknowledge prior work. Although their primary function is to inform and redirect the reader, citations are also frequently used for other purposes, such as serving as proxies of scientific impact in various types of analysis (Kanellos et al., 2021). In such cases, understanding the exact intent of a citation is crucial, as not all types of citations should be considered. For instance, while measuring the impact of an article, citations that criticize the work should generally not be considered as contributing to its impact.
Predicting citation intent based on its context (i.e., the sentences in the manuscript that accompany the citation) and other related information, has become an important classification problem to support the aforementioned use cases. Existing approaches have traditionally relied on linguistic features (Jurgens et al., 2018), machine learning methods (Teufel et al., 2006), and, in recent years, on domain-specific pre-trained language models (PLMs), such as SciBERT (Beltagy et al., 2019), that require large scientific datasets such as (Lo et al., 2020) and task-specific architectures111In this paper, when referring to traditional PLMs in this context, we primarily mean models based on the encoder-only Transformer architecture, such as BERT and its derivatives (e.g., SciBERT), which were commonly fine-tuned for specific classification tasks. This distinguishes them from the general-purpose Large Language Models (LLMs) evaluated in this work, which typically utilize decoder-only or encoder-decoder Transformer architectures and are often leveraged for their generative and in-context learning capabilities.. In contrast, this work is the first to explore the potential of open, general-purpose Large Language Models (LLMs) – which, despite potentially encountering scientific text during their broad pre-training, are not specifically optimized for the scientific domain like SciBERT – to accurately identify citation intent, evaluating their effectiveness through in-context learning and fine-tuning on minimal task-specific data.
@@
In-context learning is a paradigm that allows LLMs to learn tasks given only a few examples in the form of demonstration (Brown et al., 2020; Dong et al., 2024). In-context learning is particularly suited to citation intent classification, as it allows the models to leverage their contextual understanding of language to make accurate predictions.
@@ 1.1 Models
For our experiments, we selected the Instruction-tuned versions of five prominent open-weight model families: LLaMA (Grattafiori et al., 2024), Mistral (Jiang et al., 2023), Phi (Abdin et al., 2024), Gemma (Gemma et al., 2024), and Qwen (Bai et al., 2023; Qwen et al., 2025). A key challenge was performing the experiments on commodity hardware, both due to computational limitations and to demonstrate that citation intent classification can be executed efficiently with limited resources. This influenced the selection of models, which range in size from small (B parameters) to medium-sized (B parameters).
@@
To reduce the memory footprint and computational requirements of our evaluation, we utilize the 8-bit (Q8) quantized versions of the models. This approach significantly reduces memory usage without compromising performance. Quantization involves converting model parameters from higher numerical precisions (e.g. 16-bit floating-point) to formats of lower numerical precision (e.g. 8-bit integers). This process enables more efficient computation and reduced memory usage while largely preserving model performance and expressive power (Jurafsky and Martin, 2025).
@@ 1.2 Datasets
SciCite (Cohan et al., 2019) consists of citation strings annotated with three classes: Background Information, Method, and Results Comparison.
ACL-ARC (Jurgens et al., 2018) contains citation strings split into six categories:
Background, Motivation, Uses, Extends, Compares or Contrasts, and Future.
@@ 1.3 Configuration Parameters
To evaluate model performance, we applied four prompting methods444Throughout the paper, we use “In-context Learning Methods” and “Prompting Methods” interchangibly. by following the In-Context Learning paradigm (Brown et al., 2020): Zero-shot (no examples), One-shot (a single example per class), Few-shot ( examples per class), and Many-shot ( examples per class).
These prompting methods were selected to evaluate whether model performance improves as the number of examples increases, and to identify any saturation point where additional examples yield diminishing or negative returns. This aligns with prior literature on in-context learning methods, where these specific configurations have been extensively studied (Dong et al., 2024). Incorporating examples directly into the prompts allows models to better understand the task and the expected output format, which is especially important for citation intent classification, where the citation context is key to determining the correct label.
@@
SP2: This prompt introduced more structure, drawing inspiration from the CO-STAR framework (Teo, 2023). We adapted relevant CO-STAR components by definint the task Context, the Objective of the model, and detailed Response Rules. Given that our task requires a single, specific class label output, the stylistic components of CO-STAR (Style, Tone, and Audience) were not incorporated.
@@
Inline: Example citation sentences and their corresponding classes (structured according to the defined query templates) are embedded directly and sequentially within the main prompt, immediately preceding the actual query sentence that the model must classify.
@@
Temperature is a hyperparameter that controls the randomness or creativity of a language model’s outputs by adjusting the probability distribution of possible next tokens (Jurafsky and Martin, 2025). Lower temperatures (i.e., close to ) correspond to greedy decoding, where the model deterministically selects the most probable token at each step, while higher temperatures (closer to ) introduce greater variability by allowing the model to sample from a broader range of options.
@@ 2 Experimental Results and Analysis
This section presents the results of our experiments, focusing on model performance across various configurations and parameter settings. An initial overview of the peak F1-score achieved by each of the twelve models on both the SciCite and ACL-ARC datasets is summarized in Tables 1 and 2. The subsections that follow provide a detailed analysis of top-performing model configurations, evaluate overall model rankings, and explore the influence of specific parameters.
Rank | Model | F1-Score |
1 | Qwen 2.5 – 14B | 78.33 |
2 | Gemma 2 – 27B | 77.86 |
3 | Mistral Nemo – 12B | 77.39 |
4 | Gemma 2 – 9B | 75.12 |
5 | Phi 3 Medium – 14B | 74.67 |
6 | LLaMA 3 – 8B | 74.39 |
7 | Qwen 2 – 7B | 72.89 |
8 | LLaMA 3.1 – 8B | 72.46 |
9 | Gemma 2 - 2B | 68.79 |
10 | Phi 3.5 Mini – 3.8B | 68.25 |
11 | LLaMA 3.2 - 3B | 67.99 |
12 | LLaMA 3.2 - 1B | 45.44 |
@@ 2.1 Model Performance over Configurations
Rank | Model | F1-Score |
1 | Qwen 2.5 – 14B | 63.68 |
2 | Gemma 2 – 27B | 58.95 |
3 | Gemma 2 – 9B | 57.19 |
4 | Qwen 2 – 7B | 51.26 |
5 | LLaMA 3.1 – 8B | 48.45 |
6 | Mistral Nemo – 12B | 48.11 |
7 | Phi 3.5 Mini – 3.8B | 43.74 |
8 | Phi 3 Medium - 14B | 43.46 |
9 | Gemma 2 - 2B | 40.96 |
10 | LLaMA 3.2 - 3B | 40.07 |
11 | LLaMA 3 - 8B | 38.06 |
12 | LLaMA 3.2 - 1B | 24.60 |
@@
2.1.2 Ranked Evaluation
While the Best-Performer Evaluation provided insights into the most dominant models across configurations, it lacked the ability to capture nuances among high-performing models. For instance, a model that consistently ranked second in multiple configurations would not be reflected in that approach. To address this limitation, we adopted a ranking methodology inspired by Reciprocal Rank Fusion (RRF) (Cormack et al., 2009). This approach allowed us to evaluate all models considering their relative performance across different configurations.
@@ 2.1.2 Ranked Evaluation
Tables 4 and 5 present the results of this ranked evaluation. The ranked evaluation aligns with the Best-Performer Evaluation in identifying Qwen 2.5 14B as the most dominant model. This view also allows us to gain insights into the relative performance of other the models, highlighting distinctions among high performers, as well as the performance of lower-ranked models.
2.2 Parameter Performance Analysis
Rank | Model | Ranking Score |
1 | Qwen 2.5 – 14B | 144.163 |
2 | Mistral Nemo – 12B | 68.479 |
3 | Gemma 2 – 27B | 63.369 |
4 | Gemma 2 – 9B | 59.760 |
5 | Qwen 2 – 7B | 38.617 |
6 | LLaMA 3.1 – 8B | 27.624 |
7 | LLaMA 3 – 8B | 26.159 |
8 | Phi 3.5 Mini – 3.8B | 22.088 |
9 | Phi 3 Medium – 14B | 19.370 |
10 | Gemma 2 - 2B | 19.253 |
11 | LLaMA 3.2 - 3B | 18.286 |
12 | LLaMA 3.2 - 1B | 14.136 |
Rank | Model | Ranking Score |
1 | Qwen 2.5 – 14B | 146.496 |
2 | Gemma 2 – 27B | 62.973 |
3 | Gemma 2 – 9B | 55.318 |
4 | Qwen 2 – 7B | 41.496 |
5 | LLaMA 3.1 – 8B | 40.844 |
6 | Phi 3.5 Mini – 3.8B | 30.470 |
7 | Mistral Nemo – 12B | 30.285 |
8 | LLaMA 3.2 - 3B | 23.984 |
9 | Phi 3 Medium - 14B | 21.269 |
10 | LLaMA 3.2 - 1B | 16.796 |
11 | Gemma 2 - 2B | 16.534 |
12 | LLaMA 3 - 8B | 16.359 |
@@ 2.2 Parameter Performance Analysis
2.2.1 Chi-Square Statistical Test
To complement our parameter performance analysis, we also conducted a Chi-Square Test of Independence (Greenwood and Nikulin, 1996) to evaluate the relationship between parameter settings and F1-scores. This approach allowed us to determine whether specific parameter settings were significantly associated with performance outcomes and validate the results of our primary analysis. The results of the test (summarized in Table 7) largely reinforces the findings from our quantile-based analysis regarding the influence of different parameter settings.
@@ 2.2.1 Chi-Square Statistical Test
3 Fine-tuning
3.1 Training Configuration
For this experiment, we used the SciCite and ACL-ARC datasets, which were converted into the Alpaca format (Taori et al., 2023). This format includes a system prompt, an instruction, the citing sentence, and the true label. The use of Supervised Fine-Tuning (SFT) was motivated by its ability to adapt pre-trained models to specific tasks using minimal labeled data, making it an effective approach for citation intent classification. To perform the fine-tuning, we used the original training set of each dataset (8,243 examples for SciCite and 1,688 for ACL-ARC). For testing, we used the original test sets as seen in the previous experiments.
To fine-tune the Qwen 2.5 14B Instruct model, we used LLaMA-Factory (Zheng et al., 2024) on an AWS g6e.12xlarge EC2 instance equipped with 4 NVIDIA L40S GPUs, providing a combined GPU memory of 192GB. The training process was conducted using fp16 mixed precision to optimize memory usage and speed. To prevent memory issues during fine-tuning, we employed DeepSpeed ZeRO Stage 3 Offload (Ren et al., 2021), which enabled efficient memory management by offloading optimizer states and gradients to the CPU. The training parameters included a learning rate of 5e-5, a batch size of 16, and 10 epochs. The model was optimized using AdamW, with a cutoff length of 512 tokens for input sequences.
To optimize the fine-tuning process, we employed Low-Rank Adaptation (LoRA), a widely used parameter-efficient fine-tuning (PEFT) method (Hu et al., 2021; Han et al., 2024). LoRA enables efficient adaptation of large language models by freezing the pre-trained model weights and introducing trainable low-rank matrices into the Transformer layers, significantly reducing the number of trainable parameters while maintaining performance. For this experiment, we configured LoRA with a rank of 8, alpha of 16, and a dropout of 0.1, which allowed us to fine-tune the model effectively without exceeding memory constraints. The outlined training parameters are also summarized in Table 8.
Model | SciCite | ACL-ARC |
Instruct Q8 | 78.33 | 63.68 |
Instruct FP16 | 80.41 | 65.64 |
Fine-tuned Q8 | 86.47 | 68.48 |
Fine-tuned FP16 | 86.84 | 67.73 |
@@ 3.3 Discussion
We consider a key advantage of LLMs to be their ease of use and deployment. Tools such as LM Studio and Ollama888https://ollama.com/ allow models like those outlined in this paper to be deployed locally with zero technical expertise, making them accessible to users without a computer science background. In addition to their accessibility, LLMs offer significant adaptability. Unlike traditional methods, which often require complex pretraining or domain-specific tuning, LLMs can be fine-tuned for a wide range of scientometric tasks, such as citation recommendation, paper summarization, or trend analysis, without the need for bespoke architectures. Furthermore, LLMs can scale effectively to new scientific domains or datasets, requiring only small amounts of task-specific data to adapt to underexplored scientific fields; our fine-tuned models required only the several thousand citing sentences provided by our task-specific datasets for adaptation. This contrasts sharply with models like SciBERT, whose effectiveness stems from deliberate pre-training exclusively on a large corpus of scientific papers (millions of articles), optimizing it for scientific language understanding. While the general-purpose LLMs used in our study were likely exposed to scientific text within their vast, diverse pre-training data drawn from the web, this exposure is incidental rather than targeted. They were not specifically pre-trained or architected with the primary goal of processing scientific literature, unlike SciBERT. Our results demonstrate that even without such specialized scientific pre-training, general-purpose LLMs can achieve competitive performance through efficient fine-tuning on minimal task-specific data.
The promising results observed in this study suggest that the performance of LLMs in citation intent classification can be further improved with techniques such as chain-of-thought prompting and reasoning-focused models. These methods could enhance the models’ ability to better distinguish subtle differences in intents, refining their predictions and improving overall classification accuracy.